### v1

# Private Sketches for Linear Regression

Abstract
Linear regression is frequently applied in a variety of domains. In order to improve the efficiency of these methods, various methods have been developed that compute summaries or sketches of the datasets. Certain domains, however, contain sensitive data which necessitates that the application of these statistical methods does not reveal private information. Differentially private (DP) linear regression methods have been developed for mitigating this problem. These techniques typically involve estimating a noisy version of the parameter vector. Instead, we propose releasing private sketches of the datasets. We present differentially private sketches for the problems of least squares regression, as well as least absolute deviations regression. The availability of these private sketches facilitates the application of commonly available solvers for regression, without the risk of privacy leakage.
1 Introduction
The problem of differentially private (DP) linear regression aims to compute the optimal regressor privately. Given a data matrix and a response vector the linear regression problem is to compute where denotes (or least absolute deviations) regression and denotes (least squares) regression. DP linear regression algorithms estimate privately. Most of the literature on DP linear regression focusses on regression, while Wang & Xu (2022); Liu et al. (2024) proposed algorithms for DP norm regression.
Since the estimator for regression has a closed form expression, existing approaches look at combining private estimates of and Wang (2018), using random projections for the estimation Sheffet (2017), estimating the solution using private gradient descent type algorithms Varshney et al. (2022), as well as exponential mechanism type approaches Liu et al. (2021). Brown et al. (2024) proposes estimating the covariance matrix on a “good" set of data and perturbing the least square estimate. For the case of regression, Liu et al. (2024) proposes an iterative algorithm for estimating the solution which includes a kernel density estimation step for the purpose of introducing noise. Wang & Xu (2022) proposes using the exponential mechanism for computing the solution privately.
Each of these methods, except Sheffet (2017) involve multiple steps for estimating some property of the data, which is then used for making the outputs private. In this paper, we propose a different technique for the private linear regression problem. We consider the classic sketch-and-solve paradigm (Woodruff, 2014). In this paradigm, we compute a sketch of the data and then compute the solution to the regression problem on this sketch. Various sketching algorithms exist that are able to given a good approximation for both and regression problems Munteanu et al. (2023); Woodruff (2014); Nelson & Nguyên (2013); Meng & Mahoney (2013). While sketching for tasks such as private low rank approximation Upadhyay (2017) and for distributed private data analysis Burkhardt et al. (2025), to the best of our knowledge, this is the first work that has looked at releasing private sketches for and regression.
The advantage of releasing private sketches is that the regression problem can be solved using readily available solvers for both and regression instead of having to estimate various properties of the data. Moreover, these private sketches can be queried infinitely, without at loss of privacy. The idea of releasing private sketches is closely related to the idea of releasing private coresets, which has been done for clustering (Feldman et al., 2009; 2017). In this paper, we have not looked at private coresets for linear regression since coresets for regression generally depend on importance sampling from a data dependent distribution and are sensitive to changes in the dataset.
One of the most widely used sketching algorithms for regression is the Johnson Lindenstrauss transform (JLT) (Johnson & Lindenstrauss, 1984; Sarlos, 2006). In Sheffet (2017), it was shown that JLT itself preserves differential privacy when the dataset satisfies certain conditions. We look at releasing private sketches using the JLT. Sheffet (2019) also noted that adding noise to the data matrix for the purpose of private linear regression results in a ridge regression problem. The regularization coefficient in this problem controls the privacy of the obtained solution. In our proposed sketches, we also observe that adding noise to create private sketches results in having to solve sketched versions of regularized regression, both for and When using the JLT for releasing sketches, we show that the noise addition method can be altered to give a sketched version of unregularized regression problem. This has the benefit that the solution obtained on the sketch is the same that would obtained on the full data. When the sketching matrix is CountSketch or the sketching algorithm of Munteanu et al. (2023), we still have to deal with a regularized regression problem. We give bounds on the regularization coefficients in these cases.
Organization of the paper
2 Preliminaries
We consider datasets of the form where is the data matrix and is the response vector. In most of the discussions, we take where is appended to the columns of resulting in the matrix Two datasets are neighbouring if they differ in a single row. We assume that the norm of the rows of are bounded by Also, the matrix is assumed to have full column rank.
Let be the singular value decomposition (SVD) of We use the notation to denote the minimum singular value of The regression loss is denoted by where where The symbols denotes approximation error and failure probability respectively, such that We denote the indicator function by The absolute value of a quantity is denoted by The multivariate normal distribution with mean and variance will be denoted by where is the identity matrix.
We define some of the concepts that will be used in the paper.
Definition 1 (Differential Privacy Dwork et al. (2006)).
A randomized algorithm is -differentially private if for all neighbouring datasets and for all possible outputs in the range of
Definition 2 ( sensitivity).
The sensitivity of a function is defined as
Definition 3 (Gaussian Mechanism Dwork et al. (2006)).
The Gaussian mechanism with noise level is defined as
Definition 4 (The Johnson-Lindenstrauss Lemma Johnson & Lindenstrauss (1984); Sarlos (2006)).
Let and such that are independent random variables. If then for a -element subset with probability at least we have
for al The matrix is the JL Transform.
We state a result from Sarlos (2006) that we will be using in the discussions later.
Lemma 1 (Theorem Sarlos (2006)).
Let and be a JL transform. Let and Then, with probability at least
Definition 5 (Subspace Embedding Clarkson & Woodruff (2017)).
A matrix is said to be a subspace embedding for for a fixed matrix if for all
with probability at least
We will be using the result of Clarkson & Woodruff (2017) regarding approximation guarantee to the regression problem.
Lemma 2 (Theorem , Clarkson & Woodruff (2017)).
Let be a subspace embedding for the norm. Let and Then, with probability at least we have
which can be computed in
For the norm, we will be using the notion of lopsided embedding which we use here.
Definition 6 (Weak Weighted -sketch Munteanu et al. (2023)).
For the dataset and a weak weighted -sketch has together with a weight vector such that for all we have
and for we have
These type of embeddings are referred to as lopsided embeddings as contraction and dilation guarantees are different. Similar embedding guarantees also hold for sketching with the Cauchy transform (Sohler & Woodruff, 2011).
3 Private Sketching for regression
The Johnson Lindenstrauss transform (JLT) Johnson & Lindenstrauss (1984) has been shown to preserve -differential privacy when used for publishing a sanitized covariance matrix Blocki et al. (2012) of the data matrix A necessary condition for the privacy to be preserved is that must be sufficiently large. Following this, Sheffet (2019; 2017) used the JLT for estimating the private JL projection of Since the JLT is known to give a good approximation to linear regression problems Sarlos (2006), the private JL projection can be used for computing the solution.
The method proposed by Sheffet (2017) required that the norm of each row is bounded by Additionally, for a threshold such that the sketches can be computed in a privately. In case a matrix is appended to to get such that the The effect of concatenating is to shift the singular values by such that becomes greater than or equal to Let be the SVD of Then,
When let be the JLT (for a suitably chosen ). The private sketch released in this case is Let Then, from Theorem of Sarlos (2006), with constant probability, we have
In case let such that and for a suitable chosen The private sketch released is The least squares problem now becomes As noted in Sheffet (2019), i.e, it is an instance of the ridge regression problem where the parameter helps to preserve -differential privacy.
Then, we get by the right-hand inequality of the JL guarantee. From Sarlos (2006) and the JL guarantee, for we have
with constant probability. In this case, we have an approximate solution to the ridge regression problem where the regularization coefficient is As we will observe in the following discussions, releasing differentially private sketches results in having to solve a sketched ridge regression problem. We will also show the estimates of the regularization coefficients. If the privacy budget is very small, the coefficients have larger values implying that the utility of the solution is very bad.
3.1 Private JL Sketch with no regularization
Here, we show that when for and is close to (say ), then the sketching algorithm of Sheffet (2017) can be modified such that solving the regression problem on the private sketch does not require solving a regularized least squares problem. In fact, solving the regression problem on the private sketch gives the same solution as that obtained when solving the problem on a non-private sketch, though there is a slight degradation in the error when
We want to form a matrix such that Let be a matrix such that, for some scalar we append to get Similar to Equation 4, we want that concatenating increases the singular values of the resulting matrix Thus, we have If then we get This shifts the singular values of by Thus, we set and Also, note that
Let where and The private sketch released in this case is The least-squares problem on the released sketch then becomes Let Then, using the guarantees of the JLT and Theorem of Sarlos (2006), with constant probability, we have
where the equality in the fourth step comes from Thus, by concatenating with the singular values are scaled up by a factor of Solving the regression problem on the private sketch gives the solution as when solving the regression problem on the non-private sketch as well as a multiplicative error guarantee with to the minimum error on the original dataset.
Algorithm 1 presents the proposed method. The values of the input parameters are same as in Sheffet (2017). We restate the main theorem and a corollary from Sheffet (2019), which gives a bound on the sketching dimension The proof of privacy of Algorithm 1 can be found in (Sheffet, 2017).
Theorem 1 (Theorem 3.1, Sheffet (2019)).
Let where and is the bound on the norm of each row of Given that the algorithm that picks a matrix whose entries are i.i.d samples from a normal distribution and publishes is -differentially private.
3.2 Private CountSketch for regression
Here, we look at another type of sketching method, using the CountSketch matrix, and discuss the alterations such that the sketching algorithm becomes differentially private. The CountSketch matrix gives low-distortion subspace embeddings for the norm in a data oblivious manner, and has been shown to give -approximate solution to the regression problem (Clarkson & Woodruff, 2017; Nelson & Nguyên, 2013; Meng & Mahoney, 2013). In Zhao et al. (2022), it was shown that a noisy initialization (calibrated to the sensitivity) results in a differentially private count sketch for the purpose of frequency estimation. Releasing private sketches for various downstream problems have also been proposed in (Li et al., 2019; Coleman & Shrivastava, 2021).
Let be a count sketch matrix. Each column of has a single non-zero entry the coordinate for which is chosen uniformly at random. Each non-zero entry is either with probability or with probability The th row of the sketch, sums up the rows of (multiplied by ) that are present at the indexes corresponding to the non-zeros of and every row of is chosen exactly once. Thus, Let be a neighbouring dataset that does not agree with on the th row. Then, the sensitivity of computing can be bounded as
So, if we add a noise vector to each row of then the algorithm releases as the private sketch and it is -differentially private owing to the Gaussian mechanism. Here, is a matrix where each row is the noise vector added to each row for Then, the regression problem becomes But we cannot use Theorem of Sarlos (2006) to give approximation guarantees for Hence, we introduce the Gaussian noise for preserving privacy differently.
Let be a matrix such that each row of is sampled independently from We concatenate to the rows of in order to get which is a matrix. Let the countsketch matrix be where and are also countsketch matrices. Then, the private sketch to be released is The effect of is to add noise vectors to each of the rows of Now, we specify the number of rows of
In order to make the sketch private, we want to add the noise vectors from to each of the rows of Thus, we want that maps the rows of to each row of at least once. This is analogous to the Coupon Collector’s problem in that we have to collect coupons at least once in order to win. The expected number of trials required is then Thus, the noise matrix is a matrix.
Let Then, from the subspace embedding guarantees of the CountSketch matrix as well as regression approximation guarantees from Meng & Mahoney (2013), with constant probability we have
From the inequality at 9, we observer that solving the least squares regression problem on the private sketch results in solving a sketched ridge regression problem. In Lemma 3, we bound the regularization coefficient.
Lemma 3.
Let be the sketching matrix which has been described above and let Also, let be a matrix where the th row Then, solving the regression problem on is same as solving a sketched ridge regression problem where the regularization coefficient is
with constant probability.
Proof.
Let Each is sampled from So, we have We are interested in bounding where is an -dimensional vector of normal random variables. From Ledoux & Talagrand (1991), for we have
Here, Suppose with probability at most Then,
4 Private Sketching for regression
An analogue of the JLT for the problem of -norm is the Cauchy transform (Sohler & Woodruff, 2011; Clarkson et al., 2016), though it has weaker guarantees. The Cauchy transform is used for speeding up the computation of a distribution that is used for sampling from the rows of This algorithm was proposed in Clarkson (2005) to obtain a approximation to the solution of the regression problem. However, this algorithm outputs a weighted subsample from the dataset itself, which could lead to the leakage of private information.
Instead we look at a recently proposed oblivious linear sketching technique that satisfies the conditions of a lopsided embedding and gives an approximation to the solution of the regression problem (Munteanu et al., 2023). Similar sketches have also been utilized in (Clarkson & Woodruff, 2015; Munteanu et al., 2021). The sketch is analogous to a concatenation of multiple CountMinSketches. Similar to Section 3.2, we append a noise matrix to the rows of and release a sketch of this augmented matrix as the private sketch. We use the sketching algorithm of Munteanu et al. (2023).
4.1 An Illustration
For the purpose of illustration, we consider a simpler sketching matrix where each column of has a single entry set to Also, for any vector assume that with constant probability. The th row of can be written as Let be the neighbouring dataset such that and differ in the th row. Then, the sensitivity of can be bounded as
Thus, if we add a noise vector then the algorithm that samples and releases is -DP, using the Gaussian mechanism. But, as discussed in Section 3.2, we will not be able to give approximation guarantees on the sketched problem , Here, is a noise matrix for which each row is
Instead we append a noise matrix to the rows of to get and release the sketched matrix as the private sketch. Using the Coupon Collector’s argument as before, is a matrix. Solving the regression problem on is same as solving a sketched regularized regression problem. Let (Gordon et al., 2006). Then, with constant probability, we have
We will be using Lemma 4 for showing a bound on the regularization coefficient.
Lemma 4.
Let be a vector of independent and identically distributed (i.i.d) random variables with Let Then, with probability at least
Proof.
Since the random variable has a half normal distribution. For the moment generating function (MGF) of is
using the property that Since, s are i.i.d random variables, so too are the s. Then, for , the MGF of becomes Applying Chernoff bound on the sum of i.i.d random variables, we have
Let Then, we get
since The minimizer of is at At the minimizer, we get which gives Thus,
If then
This leads us to Lemma 5.
Lemma 5.
With constant probability,
Proof.
We have where is the th row of Also, each coordinate of is sampled i.i.d from Thus, Therefore, is the sum of i.i.d random variables each sampled from Hence, by Lemma 4, we have
with probability at least ∎
Now, we consider the sketching algorithm of Munteanu et al. (2023) and apply the modifications to make it private.
4.2 Private Sketch
The sketching algorithm given in Munteanu et al. (2023) outputs a weak weighted sketch that obtains an approximation to the regression problem. We restate the guarantee from (Munteanu et al., 2023).
Theorem 2 (Theorem 1, Munteanu et al. (2023)).
For there is a distribution over sketching matrices and a corresponding weight vector for which can be computed in time in a single pass over a turnstile data stream such that is a weak weighted -sketch for -regression with failure probability at most where
for any constant and and are constant,
and
Since, we append a noise matrix to (to get the noise augmented matrix ), the weight vector is also of length Note that the weight vector as well as the sketch is data oblivious. We give a brief description of the sketching algorithm in Munteanu et al. (2023). The sketching matrix consists of levels of sketch as shown.
Here, is the number of sketch levels, where is a branching parameter. The number of buckets(rows) at level is denoted by The probability that any row is sampled at level is The rows of that are mapped to the same row of the sketch are added up. Also, the weight of any bucket at level is set to
The sampling probabilities decrease exponentially as the levels increase. At level sketching with is a CountMinSketch of the entire dataset. It captures the elements that make a significant contribution to the objective, the so-called heavy hitters. For improving efficiency, the rows at level are mapped to rows of the sketch. On the other hand, the sketch at level corresponds to a small uniform subsample of the data. For each sketch level the probability of having a at any of the columns is For a detailed analysis of the algorithm, readers may refer to (Munteanu et al., 2023).
The private sketch, sums up the corresponding rows of and at least one of the noisy rows from The details are given in Algorithm 2. It is similar to the algorithm of Munteanu et al. (2023) with a few alterations.
In Algorithm 2, a row can be sampled at most times. Therefore, the sensitivity of the applying the sketch is at most We can now state Theorem 3.
Theorem 3.
Proof.
The differential privacy guarantee comes from the application of the Gaussian mechanism. The sensitivity of the sketch is Applying Lemma 5, we get the bound on the additive error. ∎
5 Conclusion and Future Work
In this paper, we proposed a few techniques for releasing differentially private sketches of the dataset. We utilized the JL transform as well as sparse sketches for computing the sketches. We observed that in most of the cases, introducing noise in order to preserve privacy results in having to solve sketch regularized regression problems. The bounds on the regularization coefficients highlight the regularization effect of maintaining privacy. The coefficients could potentially be large, which could lead to a large drop in utility of the solutions on the private sketches. For the JL transform, we showed that it is possible to get unregularized regression formulations in certain cases. A future direction of work is to come up with similar constructions for the other problems, especially regression, as well.