### v2

# Private Sketches for Linear Regression

Abstract
Linear regression is frequently applied in a variety of domains, some of which might contain sensitive information. This necessitates that the application of these methods does not reveal private information. Differentially private (DP) linear regression methods, developed for this purpose, compute private estimates of the solution. These techniques typically involve computing a noisy version of the solution vector. Instead, we propose releasing private sketches of the datasets, which can then be used to compute an approximate solution to the regression problem. This is motivated by the sketch-and-solve paradigm, where the regression problem is solved on a smaller sketch of the dataset instead of on the original problem space. The solution obtained on the sketch can also be shown to have good approximation guarantees to the original problem. Various sketching methods have been developed for improving the computational efficiency of linear regression problems under this paradigm. We adopt this paradigm for the purpose of releasing private sketches of the data. We construct differentially private sketches for the problems of least squares regression, as well as least absolute deviations regression. We show that the privacy constraints lead to sketched versions of regularized regression. We compute the bounds on the regularization parameter required for guaranteeing privacy. The availability of these private sketches facilitates the application of commonly available solvers for regression, without the risk of privacy leakage.
1 Introduction
Differentially private (DP) linear regression methods aim to compute the solution to the problem privately. Given a data matrix and a response vector the linear regression problem is to compute where denotes (least absolute deviations) regression and denotes (least squares) regression. DP linear regression algorithms estimate the solution privately. Most of the literature on DP linear regression focusses on regression, while Wang & Xu (2022); Liu et al. (2024) proposed algorithms for DP norm regression.
Since the solution for regression has a closed form expression, existing approaches include computing private estimates of and Wang (2018), using random projections for the privately estimating Sheffet (2017), estimating the solution using private gradient descent type algorithms Varshney et al. (2022), as well as exponential mechanism type approaches Liu et al. (2021). Brown et al. (2024) proposes estimating the covariance matrix on a “good” set of data (discarding points with high leverage scores as well as high residuals) and perturbing the least squares estimate. For the case of regression, Liu et al. (2024) proposes an iterative algorithm for estimating the solution which includes a kernel density estimation step for the purpose of introducing noise. Wang & Xu (2022) proposes using the exponential mechanism for computing the solution privately.
Each of these methods, except Sheffet (2017), involve multiple steps for estimating some property of the data matrix, which is then used for making the solutions private. In this paper, we propose releasing differentially private sketches of the data for solving the linear regression problem. We consider the classic sketch-and-solve paradigm (Woodruff, 2014) where an approximate solution to the regression problem is computed on a sketch of the data matrix. The sketched data matrix has much fewer rows than the original data matrix due to which computing the solution on the sketch is computationally efficient. Various sketching algorithms exist that are able to given a good approximation for both and regression problems Sarlos (2006); Nelson & Nguyên (2013); Meng & Mahoney (2013); Woodruff (2014); Clarkson & Woodruff (2017); Munteanu et al. (2021; 2023). While sketching has been utilized for tasks such as private low rank approximation Upadhyay (2017) and distributed private data analysis Burkhardt et al. (2025), to the best of our knowledge, this is the first work that has looked at releasing private sketches for and regression. We note that Sheffet (2017) utilizes random projections of the data matrix for computing the solution to regression privately.
An advantage of releasing private sketches of the data matrix is that the sketched regression problem can be solved using readily available solvers for both and regression, instead of having to estimate various properties of the data. Additionally, these private sketches can be queried infinitely, without any loss of privacy. The idea of releasing private sketches is closely related to the idea of releasing private coresets, which has been done for clustering (Feldman et al., 2009; 2017). In this paper, we have not looked at private coresets for linear regression since coresets for regression generally depend on importance sampling from a data dependent distribution and are sensitive to changes in the dataset.
One of the most widely used sketching algorithms for regression is the Johnson Lindenstrauss transform (JLT) (Johnson & Lindenstrauss, 1984; Sarlos, 2006). In Sheffet (2017), it was shown that JLT itself preserves differential privacy when the data matrix satisfies certain conditions. Sheffet (2019) also noted that adding noise to the data matrix for the purpose of private linear regression results in a ridge regression problem where the regularization coefficient can be set such that the solution is differentially private. Note that this is unlike the standard regularized regression problems, where the regularization coefficient is set in order to the minimize the risk. In this work, we show that when noise is added in order to make the sketches private, solving the regression problem on the released private sketch is same as solving a sketched regularized regression problem (for both and ). We derive bounds for the regularization coefficient in such cases. However, when using the JLT for releasing private sketches, we show that the noise addition method can be altered such that the regression problem on the sketch is actually a sketched version of an “unregularized” regression problem. When the sketching matrix is CountSketch or the sketching algorithm of Munteanu et al. (2023), we still have to deal with a regularized regression problem.
Organization of the paper
2 Preliminaries
We consider datasets of the form where is the data matrix and is the response vector. In most of the discussions, we take where is appended to the columns of resulting in the matrix Two datasets are neighbouring if they differ in a single row. We assume that the norm of the rows of are bounded by Also, the matrix is assumed to have full column rank.
Let be the singular value decomposition (SVD) of We use the notation to denote the minimum singular value of The regression loss is denoted by where The symbols denotes approximation error and failure probability respectively, such that We denote the indicator function by The absolute value of a quantity is denoted by The multivariate isotropic normal distribution with mean and variance will be denoted by where is the identity matrix.
We define some of the concepts that will be used in the paper.
Definition 1 (Differential Privacy Dwork et al. (2006)).
A randomized algorithm is -differentially private if for all neighbouring datasets and for all possible outputs in the range of
Definition 2 ( sensitivity).
The sensitivity of a function is defined as
Definition 3 (Gaussian Mechanism Dwork et al. (2006)).
The Gaussian mechanism with noise level is defined as
The Gaussian mechanism is - differentially private.
Definition 4 (The Johnson-Lindenstrauss Lemma Johnson & Lindenstrauss (1984); Sarlos (2006)).
Let and such that are independent random variables. If then for a -element subset with probability at least we have
for all The matrix is the JL Transform.
Definition 5 (Subspace Embedding Clarkson & Woodruff (2017)).
A matrix is said to be a subspace embedding for for a fixed matrix if for all
with probability at least
Sohler & Woodruff (2011); Woodruff & Zhang (2014) showed that subspace embeddings for the norm have dilation. The sketch and solve paradigm for linear regression Woodruff (2014) consists of two steps :
1.
Compute a sketch , of the data matrix where is the sketching matrix. The number of rows in is less than in
2.
Compute on the smaller sketch.
Previous works have shown that the solution obtained on the sketched data is a good approximation to the original problem. For regression, it was shown in Sarlos (2006); Clarkson & Woodruff (2017) that with high probability. Also, for the case of regression, Munteanu et al. (2023) constructed a sketching matrix which was able to obtain an approximation to the original problem.
3 Private Sketching for regression
The Johnson Lindenstrauss transform (JLT) Johnson & Lindenstrauss (1984) has been shown to preserve -differential privacy when used for publishing a sanitized covariance matrix Blocki et al. (2012) of the data matrix A necessary condition for the privacy to be preserved is that must be sufficiently large. Following this, Sheffet (2017; 2019) used the JLT for computing a private JL projection of This private sketch was then used for computing the solution to the regression problem.
The method proposed by Sheffet (2017) required that for a threshold such that the algorithm for computing the sketches is differentially private. In case a matrix is appended to the rows of to get such that the The effect of concatenating is to shift the singular values by such that becomes greater than or equal to (Blocki et al., 2012).
When let be the JLT (for a suitably chosen ). The private sketch released in this case is Let Then, for using the guarantees of JLT and from Theorem of Sarlos (2006), with probability at least we have
Here, since no alterations have to be made in order to guarantee privacy, solving the regression problem on is same as solving the problem on a sketch of the unregularized problem.
In case let such that and for The private sketch released is The least squares problem now becomes As noted in Sheffet (2019), i.e, it is an instance of the ridge regression problem where the parameter is set to preserve -differential privacy.
Then, we get by the JL guarantee. From Sarlos (2006) and the JL guarantee, for we have
with constant probability. In this case, since we have to append in order to guarantee privacy, solving the regression problem on is same as solving the problem on a sketch of a ridge regression problem where the regularization coefficient is (Sheffet, 2017).
3.1 Private JL Sketch with no regularization
Here, we show that when for and is close to then the sketching algorithm of Sheffet (2017) can be modified such that solving the regression problem on the private sketch does not require solving a regularized least squares problem. We want to form a matrix such that Let be a matrix such that, for some scalar we append to get Similar to Equation 3, we want that concatenating increases the singular values of the resulting matrix Thus, we have If then we get This shifts the singular values of by Thus, we set and Also, note that
Let where and for The private sketch released in this case is The least-squares problem on the released sketch then becomes Let Then, using the guarantees of the JLT and Theorem of Sarlos (2006), with constant probability, we have
where the equality in the fourth step comes from Thus, by concatenating with the singular values are scaled up by a factor of Solving the regression problem on the private sketch is same as solving an unregularized regression problem.
Algorithm 1 presents the proposed method. The values of the input parameters are same as in Sheffet (2017). The proof of privacy of Algorithm 1 is same as in (Sheffet, 2017).
3.2 Private CountSketch for regression
Here, we look at another type of sketching method, using the CountSketch matrix, and discuss the alterations such that the sketching algorithm becomes differentially private. The CountSketch matrix gives low-distortion subspace embeddings for the norm in a data oblivious manner, and has been shown to give -approximate solution to the regression problem (Clarkson & Woodruff, 2017; Nelson & Nguyên, 2013; Meng & Mahoney, 2013). In Zhao et al. (2022), it was shown that a noisy initialization (calibrated to the sensitivity) results in a differentially private count sketch for the purpose of frequency estimation. Releasing private sketches for various downstream problems have also been proposed in (Li et al., 2019; Coleman & Shrivastava, 2021).
Let be a count sketch matrix with Each column of has a single non-zero entry the coordinate for which is chosen uniformly at random. Each non-zero entry is either with probability or with probability The th row of the sketch, sums up the rows of (multiplied by ) that are present at the indexes corresponding to the non-zeros of and every row of is chosen exactly once. Thus, Let be a neighbouring dataset that does not agree with on the th row. Then, the sensitivity of computing can be bounded as
So, if we add a noise vector to each row of then the algorithm releases as the private sketch and it is -differentially private owing to the Gaussian mechanism. Here, is a matrix where each row is a noise vector Then, the regression problem becomes But we cannot use Theorem of Clarkson & Woodruff (2017) to give approximation guarantees for Hence, we introduce the Gaussian noise for preserving privacy differently.
Let be a matrix such that each row of is sampled independently from We concatenate to the rows of in order to get which is a matrix. Let the countsketch matrix be where and are also countsketch matrices. Then, the private sketch to be released is The effect of is to add noise vectors to each of the rows of Now, we specify the number of rows of
In order to make the sketch private, we want to add the noise vectors from to each of the rows of Thus, we want that maps the rows of to each row of at least once. This is analogous to the Coupon Collector’s problem in that we have to collect coupons at least once in order to win. The expected number of trials required is then Thus, the noise matrix is a matrix.
Let Then, from the subspace embedding guarantees of the CountSketch matrix as well as regression approximation guarantees from Meng & Mahoney (2013), with constant probability we have
From the inequality at 8, we observer that solving the least squares regression problem on the private sketch results in solving a sketched ridge regression problem. In Theorem 1, we bound the regularization coefficient.
Theorem 1.
Let be the sketching matrix which has been described above and let Also, let be a matrix where the th row Then, solving the regression problem on is same as solving a sketched ridge regression problem where the regularization coefficient is
with constant probability.
Proof.
Let Each is sampled from So, we have We are interested in bounding where is an -dimensional vector of i.i.d normal random variables. From Ledoux & Talagrand (1991), for we have
Here, Suppose with probability at most Then,
4 Private Sketching for regression
An analogue of the JLT for the problem of -norm is the Cauchy transform (Sohler & Woodruff, 2011; Clarkson et al., 2016), though it has weaker guarantees (lopsided subspace embedding) . The Cauchy transform is used for speeding up the computation of a distribution that is used for sampling from the rows of This algorithm was proposed in Clarkson (2005) to obtain a approximation to the solution of the regression problem. However, this algorithm outputs a weighted subsample from the dataset itself, which could lead to the leakage of private information.
Instead we look at a recently proposed oblivious linear sketching technique that satisfies the conditions of a lopsided embedding and gives an approximation to the solution of the regression problem (Munteanu et al., 2023). Similar sketches have also been utilized in (Clarkson & Woodruff, 2015; Munteanu et al., 2021). The sketch is analogous to a concatenation of multiple CountMinSketches. Similar to Section 3.2, we append a noise matrix to the rows of and release a sketch of this augmented matrix as the private sketch. We use the sketching algorithm of Munteanu et al. (2023).
4.1 An Illustration
For the purpose of illustration, we consider a simpler sketching matrix where each column of has a single entry set to Also, for any vector assume that with constant probability. The th row of can be written as Let be the neighbouring dataset such that and differ in the th row. Then, the sensitivity of can be bounded as
Thus, if we add a noise vector then the algorithm that samples and releases is -DP, by the Gaussian mechanism. But, as discussed in Section 3.2, we will not be able to give approximation guarantees on the sketched problem , Here, is a noise matrix for which each row is
Instead we append a noise matrix to the rows of to get and release the sketched matrix as the private sketch. Using the Coupon Collector’s argument as before, is a matrix. Solving the regression problem on is same as solving a sketched regularized regression problem. Let where (Gordon et al., 2006). Then, with constant probability, we have
We will be using Lemma 1 for deriving a bound on the regularization coefficient.
Lemma 1.
Let be a vector of independent and identically distributed (i.i.d) random variables with Let Then, with probability at least
Proof.
Since the random variable has a half normal distribution. For the moment generating function (MGF) of is
using the property that being the Gaussian error function. Since, s are i.i.d random variables, so too are the s. Then, for , the MGF of becomes Applying Chernoff bound on the sum of i.i.d random variables, we have
Let Then, we get
since The minimizer of is at At the minimizer, we get which gives Thus,
If then
This leads us to Lemma 2.
Lemma 2.
With constant probability,
Proof.
We have where is the th row of Also, each coordinate of is sampled i.i.d from Thus, Therefore, is the sum of i.i.d random variables each sampled from Hence, by Lemma 1, we have
with probability at least ∎
Now, we consider the sketching algorithm of Munteanu et al. (2023) and apply the modifications to make it private.
4.2 Private Sketch
The sketching algorithm given in Munteanu et al. (2023) outputs a weak weighted sketch that obtains an approximation to the regression problem. The sketching dimension is for constant success probability. The sketching algorithm of Munteanu et al. (2023) also outputs an oblivious weight vector
Since, we append a noise matrix to (to get the noise augmented matrix ), the weight vector is also of length Note that the weight vector as well as the sketch is data oblivious. We give a brief description of the sketching algorithm in Munteanu et al. (2023). The sketching matrix consists of levels of sketch as shown.
Here, is the number of sketch levels, where is a branching parameter. The number of buckets(rows) at level is denoted by The probability that any row is sampled at level is The rows of that are mapped to the same row of the sketch are added up. Also, the weight of any bucket at level is set to
The sampling probabilities decrease exponentially as the levels increase. At level sketching with is a CountMinSketch of the entire dataset. It captures the elements that make a significant contribution to the objective, the so-called heavy hitters. For improving efficiency, the rows at level are mapped to rows of the sketch. On the other hand, the sketch at level corresponds to a small uniform subsample of the data. For each sketch level the probability of having a at any of the columns is For a detailed analysis of the algorithm, readers may refer to (Munteanu et al., 2023).
The private sketch, sums up the corresponding rows of and at least one of the noisy rows from The details are given in Algorithm 2. It is similar to the algorithm of Munteanu et al. (2023) with a few alterations.
In Algorithm 2, a row can be sampled at most times. Therefore, the sensitivity of the applying the sketch is at most We can now state Theorem 2.
Theorem 2.
Algorithm 2 satisfies -differential privacy and outputs a weighted sketch Solving the regression problem on is same as solving a sketched regularized regression problem where the regularization coefficient is
with constant probability.
Proof.
The differential privacy guarantee comes from the application of the Gaussian mechanism. The sensitivity of the sketch is Applying Lemma 2, we get the bound on the regularization coefficient. ∎
5 Conclusion and Future Work
In this paper, we proposed a few techniques for releasing differentially private sketches of the dataset. We utilized the JL transform as well as sparse sketches for computing the sketches. We observed that in most of the cases, introducing noise in order to preserve privacy results in having to solve sketch regularized regression problems. The bounds on the regularization coefficients highlight the regularization effect for guaranteeing privacy. The coefficients could potentially be large, which could lead to a large drop in utility of the solutions on the private sketches. For the JL transform, we showed that it is possible to get unregularized regression formulations in certain cases. While for the regression, Algorithm 1 works, we explored the feasibility of releasing private CountSketch as well. A future direction of work is to come up with similar constructions for the other problems, especially regression, as well. Specifically, we would like to have a better noise addition mechanism such that the regularization coefficient does not become large.